OpenStudy (anonymous):

evaluate the antiderivative of x e^x^2 de from 0 to 3

OpenStudy (anonymous):

Let u=2x

OpenStudy (anonymous):

I'm sorry, u=x^2

OpenStudy (anonymous):

ok du 2x

OpenStudy (anonymous):

aswer = 1/2 e^x^2 + c

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Did you mean to say evaluate integral?

OpenStudy (anonymous):

si

OpenStudy (anonymous):

si sorry

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Your new integral\[\int\limits_{}^{}e ^{u}(du)/2\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

taking the antiderivative of that will be=e^u + c /2

OpenStudy (anonymous):

= e^x^2 + c/2

OpenStudy (anonymous):

or the same as 1/2 e^x^2 + c

OpenStudy (anonymous):

Si, now you have evaluate 0 to 3.

OpenStudy (anonymous):

ohhhhhhhhh

OpenStudy (anonymous):

So no constant it is a definite integral

OpenStudy (anonymous):

:) stuped

OpenStudy (anonymous):

what u mean with no constant is a definite integral?

OpenStudy (anonymous):

OK LISTEN VERY CAREFULLY, you need to do a thing called a guassian integral for this one, multily the first integral by integral from 0 to 3 of e^y^2 and then turn into polar coordinates and do rdrd(theta) and there you go x^2 +y^2 turns into r^2 and you have the derivitive sitting right there hope this helps and this is actually how you do the bell curve for stats. DONT FORGET TO DO THE SQUAR OF THAT BECAUSE YOU MULTIPLIED BY ANOTHER INTEGRAL

OpenStudy (anonymous):

this is like chinesse.. i am in cal 2.. maybe is a easy way before getting there?

OpenStudy (anonymous):

never heard about guassian integral yet..

OpenStudy (anonymous):

can u show me a normal way?

OpenStudy (anonymous):

I use wolfram as a check. They use the same method we do here. http://www.wolframalpha.com/input/?i=integral+x+e^x^2

OpenStudy (anonymous):

put u=x^2..du=2xdx...x-->0,u-->0,x-->3,u-->9..put these values in the integral and u will get 1/2(e^x^2)-1/2

OpenStudy (anonymous):

you should get \[\int\limits_{0}^{2\pi}\}\{\int\limits_{0}^{3}\]e^r^2(rdrd(theta) which you should get \[\sqrt{\pi(e^3-1}/2\]

OpenStudy (anonymous):

haha calc 2 you have to do the mcclarean and taylor expansion and series for this GOOD LUCK

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

hahah so sorry i didnt see the xe^x^2 this one you jjust do a simple u sub

OpenStudy (anonymous):

.. omg i was going to cry

OpenStudy (anonymous):

:) wait till calc 3 to cry you will see this

OpenStudy (anonymous):

jijiiji

OpenStudy (anonymous):

so can u show me this one... what is left

OpenStudy (anonymous):

i already integrate my answer is I= 1/2 e^x^2 + c now integarate freom 0 to 3

OpenStudy (anonymous):

You make two brackets [ ] - [ ] In the first bracket you put in the integral and plug in 3 for x; the second bracket plug in 0 for x.

OpenStudy (anonymous):

That is a subtraction sign between the two brackets.

OpenStudy (bobbyleary):

\[\int\limits_{0}^{3} x e^{x^{2}}dx\]\[\int\limits_{}^{} x e^{x^{2}}dx\]Let: u = x^2 du = 2x ---> (1/2)du = x \[1/2\int\limits_{}^{} e^{u}du = 1/2[e^u]\]Substitute u = x^2 \[1/2[e^{x^{2}}]\]and evaluate from 0 to 3. \[1/2(e^{3^{2}}-e^{0^{2}}) = (1/2)(e^9-1)\]

OpenStudy (anonymous):

how u got e^0

OpenStudy (bobbyleary):

\[e^{3^{2}} = e^9\]is evaluating at 3. And \[e^{0^{2}} = 1\]is evaluate at 0.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if u need to give the exact answer? what should it be?

OpenStudy (bobbyleary):

\[(1/2)(e^{9}-1)\]

OpenStudy (anonymous):

thanks

OpenStudy (bobbyleary):

No problem!

OpenStudy (anonymous):

Exact answer means keep the e, don't use calculator approximation.

OpenStudy (anonymous):

ohhhhh :),, good point i did not know that

OpenStudy (anonymous):

thanks guys.. understanding